OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -2, an = 3 • an-1

OpenStudy (anonymous):

The answers include -6, -18, -54, -162, -486, -1458 -2, -6, -18, -54, -162, -486 -2, -6, -3, 0, 3, 6 0, 3, -6, -3, 0, 3

OpenStudy (anonymous):

I'm going to say B because it's either B or C and C doesn't look like the right pattern.

jimthompson5910 (jim_thompson5910):

B is correct you start with -2 and you multiply each term by 3 to get the next term -2*3 = -6 -6*3 = -18 etc etc

OpenStudy (anonymous):

Ok thanks!

jimthompson5910 (jim_thompson5910):

np